OpenStudy (anonymous):

Find the y-intercept of the graph of 2x + 4y = 20.

OpenStudy (anonymous):

2x + 4y = 20. 4x=20 Y=20/4=+5 (0,+5)

hero (hero):

You did that kinda backwards

OpenStudy (anonymous):

can you help fix it

hero (hero):

To find the y-intercept, set x = 0, then solve for y: 2(0) + 4y = 20 4y = 20 y = 20/4 y = 5 (x,y) = (0,5) which means the y-intercept is 5. Alternatively, put the equation in the form y = mx + b, where b = y-intercept: 2x + 4y = 20 Subtract 2x from both sides to get 4y =-2x + 20 Divide y by 4 to get y = -.5x + 5 b = 5, therefore, the y-intercept is 5.